11. A solid gold vase of mass 20.0 kg is suspended by a rope at sea bottom. What is the Buoyant Force of the seawater on the vase? 


1
Expert's answer
2021-07-12T12:17:29-0400

The Buoyant Force is given as follows:


"F = \rho_w g V_v"

where "\rho_w =1024kg/m^2" is the density of sea water, "g = 9.8N\/kg" is the gravitational acceleration, and "V_v" is the volume of the vase.

By definition, the volume is:


"V_v = \dfrac{m}{\rho_g}"

where "\rho_g = 19300kg/m^3" is the density of gold, and "m = 20kg" is the mass of the vase.

Finally, obtain:


"F = \rho_w g\dfrac{m}{\rho_g}\\
F = 1024\cdot 9.8\cdot \dfrac{20}{19300}\approx 10.4N"

Answer. 10.4N.